# Evaluating Long-Term Memory for Long-Context Question Answering

### v1

@@ Abstract
In order for large language models to achieve true conversational continuity and benefit from experiential learning, they need memory. While research has focused on the development of complex memory systems, it remains unclear which types of memory are most effective for long-context conversational tasks. We present a systematic evaluation of memory-augmented methods using LoCoMo, a benchmark of synthetic long-context dialogues annotated for question-answering tasks that require diverse reasoning strategies. We analyse full-context prompting, semantic memory through retrieval-augmented generation and agentic memory, episodic memory through in-context learning, and procedural memory through prompt optimization. Our findings show that memory-augmented approaches reduce token usage by over 90% while maintaining competitive accuracy. Memory architecture complexity should scale with model capability, with small foundation models benefitting most from RAG, and strong instruction-tuned reasoning model gaining from episodic learning through reflections and more complex agentic semantic memory. In particular, episodic memory can help LLMs recognise the limits of their own knowledge.
Evaluating Long-Term Memory for Long-Context Question Answering
Alessandra Terranova, Björn Ross, Alexandra Birch School of Informatics, The University of Edinburgh Correspondence: a.terranova@ed.ac.uk
@@ 1 Introduction
Memory is at the core of how humans think, learn, and make decisions. Similarly, for large language model (LLM) based agents, memory is fundamental to carry out extended conversations, reason over longer time frames, learn from experience, and act in a coherent way Liu et al. (2025). LLMs often struggle to maintain coherence and accuracy in long, multi-turn interactions and very extended contexts. They also lack reliable mechanisms for metacognition, for instance they struggle with recognising the limits of their own knowledge or "knowing what they don’t know" Johnson et al. (2024).
These challenges gain more importance as context windows become larger, conversations can span hundreds of turns, and learning from experience becomes essential in real-world applications. Simply feeding models longer context is not always practical or efficient, as it increases inference cost and does not address the need for models to also learn from past information. Techniques such as retrieval-augmented generation (RAG) Lewis et al. (2020) address context limitations and allow updating and controlling data sources to improve the factuality of generated text. Additionally, recent advancements have introduced mechanisms to augment LLMs with memory components, to enhance their ability of retaining information over long contexts, and of learning from experience Shinn et al. (2023); Zhao et al. (2024).
In this paper, we examine different memory architectures, semantic (RAG and A-Mem), episodic (in-context learning with reflections), and procedural (prompt optimization), within a unified evaluation framework. We focus on long-context conversational QA, and reflect on how memory augmentation can support learning from experience and improve the models’ ability to reason about their own knowledge. Our findings show that memory-augmented approaches reduce token usage by over 90% while maintaining competitive accuracy. The complexity of the memory architecture should scale with the model’s capabilities: smaller foundation models benefit most from RAG, while more advanced instruction-tuned models from episodic memory and richer semantic memory structures. Additionally, episodic memory plays an important role in mitigating LLMs’ metacognitive limitations, helping them recognise the limits of their knowledge. This comparison establishes baselines and identifies trade-offs between performance, interpretability, and resource constraints that can inform memory system design.
@@ 2.1 Memory for LLM agents
The need for a dedicated memory component in LLM-based agents seems fundamental to capture the model’s internal reasoning, its task-specific context, historical dialogue, and evolving objectives Liu et al. (2025); Zhang et al. (2024).
Similarly to human memory, memory modules for LLMs can be categorised into long and short-term memory. Short-term memory – usually the immediate context window of the LLM – is sufficient to deal with simple tasks, but long-term memory becomes essential in more complex and realistic interaction scenarios Khosla et al. (2023). Following the organization typically used for human memory, long-term memory is organised into three different mechanisms: semantic memory stores general world knowledge and facts, making it possible for the model to provide informed responses; procedural memory encodes knowledge about how to perform the task at hand; episodic memory finally captures specific past interactions or experiences, allowing the model to gather information that can inform its future decision making Hatalis et al. (2023).
Early approaches to memory through appending conversation history to the input prompt have evolved Liu et al. (2025), leading to the use of vector embeddings for retrieving memories and selective incorporation of reasoning steps into subsequent inference calls Liu et al. (2023a). A growing body of work has been investigating dedicated memory systems for LLMs, exploring memory storage and management Zhong et al. (2024); Roucher et al. (2025). Some systems maintain logs of each agent interaction Zhong et al. (2024), while others employ read-write mechanisms that allow models to update memory content as needed Modarressi et al. (2023). Recent research has been exploring more dynamic and agent-driven approaches to memory. A-mem Xu et al. (2025) proposes a semantic memory system where each memory is structured with contextual tags and dynamically linked to related memories. This enables the system to form an evolving network of knowledge that supports updates over time and memory reorganization.
@@ 2.2 Retrieval Augmented Generation
Retrieval Augmented Generation (RAG) Lewis et al. (2020) can improve LLMs efficiency and traceability, providing a way to capture and update knowledge in a modular and more interpretable way Guu et al. (2020). Retrieval-augmented LMs can be more reliable, adaptable, and attributable thanks to the use of large-scale datastores for knowledge-intensive tasks Asai et al. (2024).
The standard RAG process involves chunking a set of documents, embedding and indexing them in a vectorstore, and retrieving relevant chunks based on semantic similarity with the query; this retrieved context can then be used to augment the LLM’s prompt for generation. Recent research has made these systems agentic by introducing more autonomy in determining what to retrieve and when, possibly having different retrievers available and refining the search strategies based on intermediate results Shao et al. (2023).
@@ 2.3 Learning through memory
A growing interest has developed in leveraging LLM-based agents’ experience to improve them without relying on costly fine-tuning or parametric updates. Traditional reinforcement learning approaches, while effective in many domains, often require extensive interaction data and prolonged training times. In response to these limitations, several frameworks have been proposed that enable language agents to learn through memory and linguistic feedback.
One such approach is Reflexion Shinn et al. (2023), which allows agents to self-improve through natural language reflection stored in an episodic memory buffer, rather than weight updates. Building on Reflexion, the Experiential Learning (ExpeL) framework Zhao et al. (2024) proposes a memory-based strategy in which agents autonomously learn from experience without gradient updates, highlighting their capacity for transfer learning and emergent reasoning abilities.
@@ 3.1 Dataset and evaluation
Our goal is to test different memory strategies on a task which models realistic long-context conversations, which is why we decided to use LoCoMo (Long-term Conversational Memory) Maharana et al. (2024). LoCoMo is a publicly available evaluation benchmark consisting of very long-context synthetic conversational data generated through a machine-human pipeline. While synthetic data has limitations regarding real-world noise and ambiguous user intent, it provides controlled evaluation conditions that allow us to isolate memory mechanism effects without confounding variables from dialogue structure or inconsistent annotation quality. While previous datasets contain dialogues with around 1,000 tokens, spanning over 4-5 sessions, LoCoMo consists of ten conversations, spread across up to 35 chat sessions, each extending over 300 turns and averaging 9,000 tokens. The dataset is granularly annotated for the question-answering and event-summarization tasks and presents a comprehensive benchmark to evaluate long-term memory in models.
An additional reason to choose this dataset, is that the questions are divided into five distinct reasoning types: single-hop, multi-hop, temporal, open-domain or world knowledge, and adversarial. In particular, temporal reasoning questions require the model to use the date of conversations, while adversarial questions do not have an answer in the given data, presenting both a challenge and possible insights into the trustworthiness of the examined systems.
For evaluation, we employ the F1 score to assess answer accuracy. As QA annotations are directly taken from the conversations as much as possible Maharana et al. (2024), we instruct the models to answer with exact words from the conversation. As a measure of relative performance, we report the average F1 ranking across categories for each approach examined, and we use the average token length per query to measure efficiency. Following Maharana et al. (2024), for adversarial questions we set the evaluation score to 1 if the generated answer contains ’no information available’ and to 0 otherwise.
Unlike previous research that used different prompts depending on the type of question Maharana et al. (2024); Xu et al. (2025), we prompt the models in the same way regardless of the question type, making the task more realistic and allowing better isolation of memory mechanism effects.
@@ 3.2 Memory for Question Answering
Short-term Long-term
@@
RAG: Retrieval-Augmented Generation
We implement RAG using a top-$k$ retrieval pipeline over the full conversation history. Our RAG component represents semantic memory, as it stores the essential knowledge that grounds an agent’s responses. At inference time, for each question, the model is provided with the top-$k$ relevant utterances and their respective timestamps, ranked by cosine similarity to the query using bge-m3 embeddings. We initially test k=5,10,20k={5,10,20} to examine the impact of context window size on performance (Appendix A), and run all of our experiments with $k=10$. The retrieved snippets are appended to the model prompt before answer generation.
Figure 1: Top figure: EpMem, episodic memory through in-context learning. The model generates answers for a sample of questions and reflects on its performance using F1 scores and labelled answers to produce natural language reflections. These previous examples and reflections are stored in episodic memory and retrieved as in-context examples to guide the model’s answers on future samples. Bottom figure: PromptOpt, procedural memory through prompt optimization. The model answers a sample of questions using an initial prompt and predictions are compared with labelled answers to compute F1 scores. A classification and optimization step is then used to update the prompt, forming procedural memory. The optimized prompt is applied to subsequent samples to guide future predictions.
@@
Procedural memory encodes how an agent should behave and respond. We start with the same prompts used for the base RAG approach (Appendix B) that define the core agent behaviour and then evolve through feedback and experience. In particular, after generating answers for all the questions related to the first conversation, we record the model’s responses, the correct labels, and its performance, and iteratively prompt the model with batches of 5 examples to use to refine its instructions, aiming to learn which approaches work best for different situations (Figure 1). This optimization follows LangMem’s implementation of procedural memory LangChain (2024) and it is achieved by a classification step, where the model is prompted to select which parts of its prompt caused errors, and an optimisation step where the model is prompted to generate an updated version of those prompt parts. We use the optimized prompt for the next sets of questions.
@@
A-Mem: Agentic Semantic Memory
We replicate the A-Mem architecture as introduced by Xu et al. (2025), where the system maintains structured memory notes representing each utterance in the conversational data. As the memory agent is presented with a new utterance for the same conversation, it updates its memory with new entries and can make decisions about updating old ones and their connections. For QA, top-10 semantic memory entries are retrieved and inserted into the prompt. We follow the authors’ implementation and prompting strategy, excluding task-specific rewording for adversarial and temporal questions.
Both RAG and A-Mem implement semantic memory, but with different levels of complexity. A-Mem supplements snippet embeddings with relevant keywords, contextual tags, and arrays of related memories. While RAG retrieves based purely on semantic similarity, A-Mem includes query expansion and provides the complete memory note context.
@@ 3.3 Implementation details
Appendix B details all the prompts we employed for our experiments. We deploy LLama 3.2-3B, Mistral-7B-v0.3, Qwen2.5-7B, and the respective instruction-tuned versions through huggingface, we access GPT-4o mini through the official OpenAI API. We report our main findings in tables 5 and3, and the full experimental results in Appendix C.
We conduct all of our experiments on a 24 GB GPU machine for a single run. For the retrieval components, we employ k=10 for snippet and A-Mem memories selection, for episodic memories we limit the number of examples to 3. We use bge-m3 text embeddings for our approaches and the all-minilm-l6-v2 model used by the authors for A-Mem. We report hyperparameter details in Appendix A.
The results on GPT-4o mini validate our findings on a widely-used commercial reasoning model. Our focus on open weights smaller models, instead, serves an important practical purpose, as these represent the scale most practitioners can deploy in resource-constrained environments.

### v2

# Metacognition in Generative AI

@@ Abstract
In order for large language models to achieve true conversational continuity and benefit from experiential learning, they need memory. While research has focused on the development of complex memory systems, it remains unclear which types of memory are most effective for long-context conversational tasks. We present a systematic evaluation of memory-augmented methods on long-context dialogues annotated for question-answering tasks that require diverse reasoning strategies. We analyse full-context prompting, semantic memory through retrieval-augmented generation and agentic memory, episodic memory through in-context learning, and procedural memory through prompt optimization. Our findings show that memory-augmented approaches reduce token usage by over 90% while maintaining competitive accuracy. Memory architecture complexity should scale with model capability, with foundation models benefitting most from RAG, and stronger instruction-tuned models gaining from episodic learning through reflections and more complex agentic semantic memory. In particular, episodic memory can help LLMs recognise the limits of their own knowledge.
@@ 1 Introduction
Memory is at the core of how humans think, learn, and make decisions. Similarly, for large language model (LLM) based agents, memory is fundamental to carry out extended conversations, reason over longer time frames, learn from experience, and act in a coherent way liu2025advances. LLMs often struggle to maintain coherence and accuracy in long, multi-turn interactions and very extended contexts. They also lack reliable mechanisms for metacognition, for instance they struggle with recognising the limits of their own knowledge or "knowing what they don’t know" johnson2024imagining.
These challenges gain more importance as context windows become larger, conversations can span hundreds of turns, and learning from experience becomes essential in real-world applications. Simply feeding models longer context is not always practical or efficient, as it increases inference cost and does not address the need for models to also learn from past information. Techniques such as retrieval-augmented generation (RAG) lewis2020retrieval address context limitations and allow updating and controlling data sources to improve the factuality of generated text. Additionally, recent advancements have introduced mechanisms to augment LLMs with memory components, to enhance their ability of retaining information over long contexts, and of learning from experience shinn2023reflexion; zhao2024expel.
In this paper, we examine different memory architectures, semantic, episodic (in-context learning with reflections), and procedural (prompt optimization), within a unified evaluation framework. We focus on long-context conversational QA, and reflect on how memory augmentation can support learning from experience and improve the models’ ability to reason about their own knowledge. Our findings show that memory-augmented approaches reduce token usage by over 90% while maintaining competitive accuracy. The complexity of the memory architecture should scale with the model’s capabilities: foundation models benefit most from RAG, while more advanced instruction-tuned models from episodic memory and richer semantic memory structures. Additionally, episodic memory plays an important role in mitigating LLMs’ metacognitive limitations, helping them recognise the limits of their knowledge. This comparison establishes baselines and identifies trade-offs between performance, interpretability, and resource constraints that can inform memory system design.
@@ 2.1 Memory for LLM agents
The need for a dedicated memory component in LLM-based agents seems fundamental to capture the model’s internal reasoning, its task-specific context, historical dialogue, and evolving objectives liu2025advances; zhang2024survey. Similarly to human memory, memory modules for LLMs can be categorised into long and short-term memory. Short-term memory – usually the immediate context window of the LLM – is sufficient to deal with simple tasks, but long-term memory becomes essential in more complex and realistic interaction scenarios khosla2023survey. Following the organization typically used for human memory, long-term memory is organised into three different mechanisms: semantic memory stores general world knowledge and facts, making it possible for the model to provide informed responses; procedural memory encodes knowledge about how to perform the task at hand; episodic memory finally captures specific past interactions or experiences, allowing the model to gather information that can inform its future decision making hatalis2023memory. Early approaches to memory through appending conversation history to the input prompt have evolved liu2025advances, leading to the use of vector embeddings for retrieving memories and selective incorporation of reasoning steps into subsequent inference calls liu2023think. A growing body of work has been investigating dedicated memory systems for LLMs, exploring memory storage and management zhong2024memorybank; smolagents. Some systems maintain logs of each agent interaction zhong2024memorybank, while others employ read-write mechanisms that allow models to update memory content as needed modarressi2023ret. Recent research has been exploring more dynamic and agent-driven approaches to memory. A-mem xu2025mem proposes a semantic memory system where each memory is structured with contextual tags and dynamically linked to related memories. This enables the system to form an evolving network of knowledge that supports updates over time and memory reorganization.
@@ 2.2 Retrieval Augmented Generation
Retrieval Augmented Generation (RAG) lewis2020retrieval can improve LLMs efficiency and traceability, providing a way to capture and update knowledge in a modular and more interpretable way Guu_Lee_Tung_Pasupat_Chang_2020. Retrieval-augmented LMs can be more reliable, adaptable, and attributable thanks to the use of large-scale datastores for knowledge-intensive tasks asai2024reliable. The standard RAG process involves chunking a set of documents, embedding and indexing them in a vectorstore, and retrieving relevant chunks based on semantic similarity with the query; this retrieved context can then be used to augment the LLM’s prompt for generation. Recent research has made these systems agentic by introducing more autonomy in determining what to retrieve and when, possibly having different retrievers available and refining the search strategies based on intermediate results shao2023enhancingretrievalaugmentedlargelanguage.
@@ 2.3 Learning through memory
A growing interest has developed in leveraging LLM-based agents’ experience to improve them without relying on costly fine-tuning or parametric updates. Traditional reinforcement learning approaches, while effective in many domains, often require extensive interaction data and prolonged training times. In response to these limitations, several frameworks have been proposed that enable language agents to learn through memory and linguistic feedback. One such approach is Reflexion shinn2023reflexion, which allows agents to self-improve through natural language reflection stored in an episodic memory buffer, rather than weight updates. Building on Reflexion, the Experiential Learning (ExpeL) framework zhao2024expel proposes a memory-based strategy in which agents autonomously learn from experience without gradient updates, highlighting their capacity for transfer learning and emergent reasoning abilities.
@@ 3.1 Dataset and evaluation
Our goal is to test different memory strategies on a task which models realistic long-context conversations, which is why we decided to use LoCoMo (Long-term Conversational Memory) maharana2024evaluating. LoCoMo is a publicly available evaluation benchmark consisting of very long-context synthetic conversational data. While synthetic data has limitations regarding real-world noise and ambiguous user intent, it provides controlled evaluation conditions that allow us to isolate memory mechanism effects without confounding variables from dialogue structure or inconsistent annotation quality. While previous datasets contain dialogues with around 1,000 tokens, spanning over 4-5 sessions, LoCoMo consists of ten conversations, spread across up to 35 chat sessions, each extending over 300 turns and averaging 9,000 tokens. The dataset is granularly annotated for question-answering requiring five distinct reasoning types: single-hop, multi-hop, temporal, open-domain or world knowledge, and adversarial. In particular, temporal reasoning questions require the model to use the date of conversations, while adversarial questions do not have an answer in the given data, presenting both a challenge and possible insights into the trustworthiness of the examined systems. Additionally, we validate our results on QMSum zhong2021qmsum, a human-annotated benchmark for query-based multi-domain meeting summarization. While the meetings are shorter and the queries do not require adversarial, temporal, or open-domain reasoning, this data can show the nuance of real-world conversations in product-design, academic, and parliamentary meetings.
For evaluation, we employ the F1 score to assess answer accuracy. As QA annotations are directly taken from the conversations as much as possible maharana2024evaluating, we instruct the models to answer with exact words from the conversation. As a measure of relative performance, we report the average F1 ranking across categories for each approach examined, and we use the average token length per query to measure efficiency. Following Maharana et al. maharana2024evaluating, for adversarial questions we set the evaluation score to 1 if the generated answer contains ’no information available’ and to 0 otherwise. Unlike previous research that used different prompts depending on the type of question maharana2024evaluating; xu2025mem, we prompt the models in the same way regardless of the question type, making the task more realistic and allowing better isolation of memory mechanism effects.
@@ 3.2 Memory for Question Answering
Short-t. Long-term
@@
RAG: Semantic Memory as Retrieval-Augmented Generation
We implement RAG using a top-$k$ retrieval pipeline over the full conversation history. Our RAG component represents semantic memory, as it stores the essential knowledge that grounds an agent’s responses. At inference time, for each question, the model is provided with the top-$k$ relevant utterances and their respective timestamps, ranked by cosine similarity to the query using bge-m3 embeddings. We initially test k=5,10,20k={5,10,20} to examine the impact of context window size on performance (Appendix A), and run all of our LoCoMo experiments with $k=10$. The retrieved snippets are appended to the model prompt before answer generation.
A-Mem: Agentic Semantic Memory
We replicate the A-Mem architecture as introduced by Xu et al. xu2025mem, where the system maintains structured memory notes representing each utterance in the conversational data. As the memory agent is presented with a new utterance for the same conversation, it updates its memory with new entries and can make decisions about updating old ones and their connections. For QA, top-10 semantic memory entries are retrieved and inserted into the prompt. We follow the authors’ implementation and prompting strategy, excluding task-specific rewording for adversarial and temporal questions. Both RAG and A-Mem implement semantic memory, but with different levels of complexity. A-Mem supplements snippet embeddings with relevant keywords, contextual tags, and arrays of related memories. While RAG retrieves based purely on semantic similarity, A-Mem includes query expansion and provides the complete memory note context.
Figure 1: Left figure: EpMem, episodic memory through in-context learning. The model generates answers for a sample of questions and reflects on its performance using F1 scores and labelled answers to produce natural language reflections. These previous examples and reflections are stored in episodic memory and retrieved as in-context examples. Right figure: PromptOpt, procedural memory through prompt optimization. The model answers a sample of questions using an initial prompt and predictions are compared with labelled answers to compute F1 scores. A classification and optimization step is then used to update the prompt, forming procedural memory.
@@
Procedural memory encodes how an agent should behave and respond. We start with the same prompts used for the base RAG approach (Appendix B) that define the core agent behaviour and then evolve through feedback and experience. In particular, after generating answers for all the questions related to the first conversation, we record the model’s responses, the correct labels, and its performance, and iteratively prompt the model with batches of 5 examples to use to refine its instructions, aiming to learn which approaches work best for different situations (Figure 1). This optimization follows LangMem’s implementation of procedural memory langmem2024 and it is achieved by a classification step, where the model is prompted to select which parts of its prompt caused errors, and an optimisation step where the model is prompted to generate an updated version of those prompt parts. We use the optimized prompt for the next sets of questions.
@@ 3.3 Implementation details
Appendix B details all the prompts we employed for our experiments. We deploy LLama 3.2-3B, Qwen2.5-7B, and the respective instruction-tuned versions through huggingface, we access GPT-4o mini through the official OpenAI API. We report our main findings in Tables 2 and4, and the full experimental results in Appendix C. We conduct all of our experiments on a 24 GB GPU machine for 5 runs, and report the mean and standard deviation. For the retrieval components, we employ k=10 for snippet and A-Mem memories selection with LoCoMo data and k=20 with QMSum. For episodic memories we limit the number of examples to 3. We use bge-m3 text embeddings for our approaches and the all-minilm-l6-v2 model used by the authors for A-Mem. We report hyperparameter details in Appendix A. The results on GPT-4o mini validate our findings on a widely-used commercial model. Our focus on open weights smaller models, instead, serves an important practical purpose, as these represent the scale most practitioners can deploy in resource-constrained environments.